INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF NICKEL FLUORIDE IN THE NEAR IN-FRARED

\underline{JAMES J. O’BRIEN}, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499; LEAH C. O’BRIEN, RACHEL A. HARRIS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The (1,0) band of the [11.1] ^{2}Π_{3/2} - X ^{2}Π_{3/2} transition of NiF has been recorded using intracavity laser absorption spectroscopy. The analysis was straight-forward and based on the known Δ_{2}F values determined from microwave spectroscopy. Results of the analysis will be presented. The gas phase NiF molecules were produced using a nickel-lined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.